# Distributed Maximum Flow in Planar Graphs

### v3

@@ 3 Preliminaries
Planar embedding. Let $G=(V,E)$ be a directed planar graph. The geometric planar embedding of $G$ is a drawing of $G$ on a plane so that edges intersect only in vertices. A combinatorial planar embedding of $G$ provides for each $v\in G$, the local clockwise order of its incident edges, such that, the ordering of edges is consistent with some geometric planar embedding of $G$. Throughout, we assume that a combinatorial embedding of $G$ is known locally for each vertex. This is done in $\tilde{O}(D)$ rounds using the planar embedding algorithm of Ghaffari and Haeupler [13].
@@
Figure 8: On the left, the primal graph $G$. On the right, the face-disjoint graph $\hat{G}$. Each vertex in $G$ has a star-center (blue) copy in $\hat{G}$, in addition to $\deg(v)$ (black) additional copies. Black edges in $\hat{G}$ are duplicates of $G$ edges, such that, connected components induced by them map to faces of $G$, e.g., the face $f_{\infty}$ of $G$ maps to the component $C_{\infty}$ in $\hat{G}$ (thick black). Dashed blue edges connect each star-center to other copies of the same vertex and solid blue edges connect two components if and only if there corresponding faces should be connected in $G^{*}$. For example, there are two such edges connecting $C_{\infty}$ to the component that maps to the face which contains $w$.
The following are the properties of $\hat{G}$ that we use. For the full definition of $\hat{G}$ and a proof of these properties see LABEL:appendix:_preliminaries_hat{G}.
@@ 4 Weighted Girth via Minor Aggregations
We provide an implementation of the minor-aggregation model in the dual graph, which immediately yields a collection of algorithmic results in the primal graph. In this section we provide an overview of the approach, where the full details and proofs appear in Appendix B.
In the Minor-Aggregation model, introduced in [42], there is a given graph $G$, where the vertices and edges of $G$ are computational units, and an algorithm works in synchronous rounds, where in each round we can either contract some of the edges, or compute an aggregate function on certain sets of disjoint vertices or edges. For the full details of the model see Appendix B. As shown in [42], the minor-aggregation model can be simulated in $\mathsf{CONGEST}$ by solving the following part-wise aggregation problem.
Definition 4.1 (Part-Wise Aggregation (PA)).
Consider a partition $\{S_{1},\ldots,S_{N}\}$ of $V$ where every $G[S_{i}]$ is connected. Assume each $v\in S_{i}$ has an $\tilde{O}(1)$-bit string $x_{v}$. The PA problem asks that each vertex of $S_{i}$ knows the aggregate function $\bigoplus_{v\in S_{i}}x_{v}$.
An aggregate operator is a function that allows to replace $\tilde{O}(1)$-bit strings by one $\tilde{O}(1)$-bit string. This is usually a simple commutative function such as taking a minimum or a sum.
Minor-Aggregation for the dual graph. Our goal is to use the face disjoint graph $\hat{G}$ to simulate a minor-aggregation algorithm on the dual graph, where the vertices of a face simulate the corresponding dual node, and each dual edge is simulated by the endpoints of the corresponding primal edge. By the above discussion, our main goal is to show how to solve the PA problem on the dual graph.

### v4

@@ 3 Preliminaries
Planar embedding. Let $G=(V,E)$ be a directed planar graph. The geometric planar embedding of $G$ is a drawing of $G$ on a plane so that edges intersect only in vertices. A combinatorial planar embedding of $G$ provides for each $v\in G$, the local clockwise order of its incident edges, such that, the ordering of edges is consistent with some geometric planar embedding of $G$. For a detailed discussion, the reader is referred to [26]. Throughout, we assume that a combinatorial embedding of $G$ is known locally for each vertex. This is done in $\tilde{O}(D)$ rounds using the planar embedding algorithm of Ghaffari and Haeupler [13].
@@
[…]
We next give the full definition of the graph $\hat{G}$. Let $\deg(v)$ denote the undirected degree of a vertex $v$ in $G$. Then, a vertex $v$ in $G$ has $\deg(v)+1$ copies in $\hat{G}$: A copy $v$ referred to as a star center (we denote by $V_{S}=V(G)$ the set of all star centers) and a copy $v_{i}$ for each local region of $v$. A local region of $v$ is defined by two consecutive edges in $v$’s local embedding. Notice that $v$ has $\deg(v)$ local regions, and that each local region belongs to a face of $G$ (it is possible that several local regions of $v$ belong to the same face). Eventually, we will have that the faces of $G$ correspond to (vertex and edge disjoint) faces of $\hat{G}$. If $i$ local regions of $v$ belong to the same face $f$ of $G$, then $v$ will have $i$ copies in the face of $\hat{G}$ that corresponds to $f$. The vertices of $\hat{G}$ are therefore $V_{S}\bigcup(\cup_{v\in G}N(v))$ where $N(v)=\{v_{1},\ldots,v_{\deg(v)}\}$.
$E_{S}$:
These edges connect every star center $v$ and its copies $v_{i}\in N(v)$.
$E_{R}$:
For each edge $(u,v)\in G$, there are two edges in $E_{R}$, one copy for each face of $G$ that contains $(u,v)$. Let $(u,v)$ be the $i$’th and $j$’th edge in the clockwise local embedding of $u$ and $v$, respectively. Then the edges added to $E_{R}$ are $(u_{i+1},v_{j})$ and $(u_{i},v_{j+1})$. This is because the face containing the $i$’th and $(i+1)$’th edges of $u$ (ii’th local region of $u$) is the face containing the $j$’th and (j1)(j-1)’th edges of $v$ ((j1)(j-1)’th local region of $v$). Symmetrically, the face containing the $i$’th and (i1)(i-1)’th edges of $u$ is the face containing the $j$’th and $(j+1)$’th edges of $v$.
$E_{C}$:
In order to simulate computations on $G^{*}$, we make sure that if two faces of $G$ share an edge $e$ (i.e. are adjacent nodes in $G^{*}$) then their corresponding faces in $\hat{G}$ are connected by a corresponding edge in $E_{C}$, such an edge connects two vertices $v_{j},v_{j+1}$ of the same star, because each of $v_{j},v_{j+1}$ corresponds to a face that contains $e$. For an edge $(u,v)$ we can add an edge in $E_{C}$ between copies of $u$ or copies of $v$, the choice of which such edge to add is arbitrary: For every edge $(u,v)$ in $G$ we add the edge $(v_{j},v_{j+1})$ to $E_{C}$ if $v$’s ID is larger than $u$’s, and $(u,v)$ is the $j$’th edge in the clockwise local embedding of $v$.
In [17], $\hat{G}$ was defined without the edges $E_{C}$, as their purposes did not require it. However, the following properties (except for Property 5) of the face-disjoint graph were proven in [17] and are yet valid. For proofs of all properties see LABEL:appendix:_preliminaries_hat{G}.
@@ 4 Weighted Girth via Minor Aggregations
We provide an implementation of the minor-aggregation model in the dual graph, which immediately yields a collection of algorithmic results in the primal graph.
4.1 Preliminaries
Definition 4.1 (Low-congestion shortcuts).
Definition 4.2 (Shortcut quality).
Definition 4.3 (Aggregation operators).
Definition 4.4 (Part-wise aggregation).
Lemma 4.5 (Part-wise aggregation via low-congestion shortcuts [14, 15, 19]).
If $G$ admits an $(\alpha,\beta)$-shortcut that can be constructed in $r$ rounds, then the PA problem can be solved on $G$ in $\tilde{O}(r+\alpha+\beta)$ rounds. In particular, the PA problem can be solved on $G$ in $\tilde{O}(r+SQ(G))$ rounds, where $r$ is the number of rounds needed for constructing low-congestion shortcuts of quality $SQ(G)$.
Since a planar graph $G$ of hop-diameter $D$ admits low-congestion shortcuts of quality $SQ(G)=\tilde{O}(D)$ [14, 15, 19] that can be constructed in $\tilde{O}(D)$ rounds we get:
Corollary 4.6.
The minor-aggregation model.
We define the minor-aggregation model as first defined by [43]. We mention that [18] has defined an extended version of the model that supports virtual nodes. However, the definitions related to the extended model are delayed until we prove that we can simulate it as well, which would simply follow from simulating the basic (original) model.
Definition 4.7 (Basic minor-aggregation model [43]).
  1. 1.
    Contraction step. Each edge $e$ chooses a value in $\{{0},{1}\}$. Contracting all edges that chose 1{1} defines a minor $G^{\prime}=(V^{\prime},E^{\prime})$ of $G$. The nodes $V^{\prime}$ of $G^{\prime}$ are subsets of $V$ called super-nodes.
  2. 2.
  3. 3.
Lemma 4.8 (Minor-aggregation model simulation in $\mathsf{CONGEST}$ via PA [43, 18]).
Any $\tau$-round minor-aggregation algorithm can be simulated in the $\mathsf{CONGEST}$ model on $G$ in $\tilde{O}(\tau\cdot N)$ rounds where $N$ is the $\mathsf{CONGEST}$ round complexity for solving the PA problem on $G$.
Recall (Lemma 4.5) that the PA problem on $G$ can be solved in $N=\tilde{O}(r+SQ(G))$ $\mathsf{CONGEST}$ rounds where $r$ is the round complexity for constructing shortcuts of quality $SQ(G)$. Therefore, the above lemma implies that any $\tau$-round minor-aggregation algorithm can be simulated in the $\mathsf{CONGEST}$ model in $\tilde{O}(\tau\cdot(r+SQ(G)))$ rounds, which is $\tilde{O}(\tau\cdot D)$ for planar graphs (Corollary 4.6).
4.2 Minor-Aggregation for the Dual Graph
4.2.1 PA on $G^{*}$ and Basic Model Simulation
Since $G^{*}$ is not in hand, a straightforward $\mathsf{CONGEST}$ simulation on $G^{*}$ might suffer a linear blow-up in the round complexity. Moreover, we cannot directly follow the method of [43] of compiling the minor-aggregation model to $\mathsf{CONGEST}$, since it is unclear how to do so on $G^{*}$. Instead, we combine their approach with part-wise aggregations on $G^{*}$ via $\hat{G}$.